# A Closer Look at TabPFN v2: Strength, Limitation, and Extension

### v1

@@ Appendix C Detailed Results
Table 8: Performance of various models on 18 large-scale datasets. The results show the mean accuracy/RMSE of different models, including MLP, Logistic Regression/Linear Regression (LR), KNN, XGBoost (XGB), Random Forest (RForest), CatBoost (CatB), ModernNCA (MNCA), RealMLP, and various versions of TabPFN v2: original TabPFN v2 (PFNv2), TabPFN v2 with K-means (PFNv2-K), TabPFN v2 with Bagging (PFNv2-B), and PFNv2* (TabPFNv2*).
Dataset MLP LR KNN XGB RForest CatB MNCA RealMLP PFNv2 PFNv2-K PFNv2-B PFNv2* PFNv2-DT
BNG(credit-a) 90.07 85.98 87.41 90.21 89.25 91.13 89.98 90.91 89.55 89.01 89.66 89.89 89.96
CDC_Indicators 86.79 86.55 86.39 86.76 86.60 86.78 86.76 86.76 86.65 86.68 86.69 86.74 86.22
Higgs 75.53 64.29 65.16 73.33 71.87 74.81 73.28 75.36 71.64 71.56 72.01 72.13 72.44
Smoking_signal 73.90 72.53 72.36 73.87 73.08 73.99 73.63 74.00 73.47 73.37 73.55 73.69 72.94
nomao 96.19 94.59 95.20 96.92 96.07 97.03 96.68 96.37 96.08 96.29 96.12 96.18 96.51
sf-police-incidents 87.84 87.84 85.87 87.68 87.84 87.87 - 87.84 87.84 87.84 87.84 87.84 87.84
Data_Crowdfunding 96.48 67.04 93.70 96.89 95.29 96.81 96.53 96.71 94.59 91.81 94.96 95.07 96.93
Fashion-MNIST 89.54 85.69 86.00 90.03 86.57 90.24 89.36 90.25 68.40 82.82 83.89 86.26 78.58
covertype 94.01 72.54 92.76 96.30 78.30 90.77 97.31 97.38 83.54 82.95 84.16 86.85 96.89
jannis 71.99 64.60 65.67 71.83 69.19 72.26 72.57 73.00 70.24 70.26 70.59 71.31 71.90
poker-hand 99.99 50.12 54.01 99.51 64.63 97.69 76.31 99.88 41.97 38.86 36.80 54.12 85.60
volkert 69.85 58.75 67.41 69.74 62.71 70.88 77.18 73.76 62.82 62.15 62.81 64.84 68.53
Airlines_DepDelay (×$\times 10^{1}$) 2.905 2.933 3.170 2.891 2.907 2.881 - 2.482 2.937 2.933 2.937 2.915 2.906
Wave_Energy_Farm (×$\times 10^{3}$) 8.199 13.19 32.29 6.917 7.294 7.173 6.148 59.05 7.214 8.375 7.063 10.506 6.489
UJIndoorLoc (×$\times 10^{0}$) 9.958 $\infty$ 9.004 10.47 23.19 9.139 5.990 65.34 66.49 7.825 7.435 9.538 7.024
blogfeedback (×$\times 10^{1}$ ) 2.387 $\infty$ 2.410 2.093 2.026 2.044 1.953 2.105 3.073 2.687 2.700 2.014 2.011
microsoft (×$\times 10^{-1}$ ) 7.577 7.782 8.284 7.514 7.566 7.453 7.573 5.077 7.735 7.981 7.720 7.612 7.952
yahoo (×$\times 10^{-1}$ ) 7.692 7.997 8.504 7.629 - 7.514 - 5.671 8.148 8.332 8.132 7.961 8.118
Table 9: Performance comparison of FT-T and MNCA with and without randomized feature tokens across tiny benchmark datasets of Ye et al. (2024b). The first two columns (FT-T and MNCA) show results using randomized tokens, while the last two columns (FT-T and MNCA) represent the original methods. All values are the mean of 15 runs with different random seeds. For classification tasks, the performance is measured by accuracy, while for regression tasks, the metric used is Root Mean Squared Error (RMSE). The table demonstrates the effect of randomized feature tokens on model performance in both types of tasks.
Dataset FT-T* MNCA* FT-T MNCA
BNG(breast-w) 98.62 98.66 98.60 98.59
BNG(cmc) 58.35 58.39 58.50 58.56
BNG(tic-tac-toe) 80.92 81.10 81.31 81.08
Cardiovascular-Disease-dataset 73.35 73.29 73.22 73.20
FOREX_audchf-day-High 52.90 72.93 51.79 75.44
FOREX_audsgd-hour-High 69.53 66.86 61.84 71.09
FOREX_cadjpy-hour-High 69.62 70.99 68.80 71.18
Gender_Gap_in_Spanish_WP 59.16 59.16 59.86 59.25
KDD 80.15 80.12 80.28 79.58
VulNoneVul 98.93 98.81 98.93 98.79
baseball 93.56 93.71 93.76 92.74
credit 75.83 75.39 75.88 72.98
dis 98.29 98.31 98.76 98.68
eye_movements_bin 59.01 61.31 59.45 90.88
jungle_chess_2pcs_raw_endgame_complete 97.11 99.67 97.61 99.50
law-school-admission-bianry 99.00 100.0 98.02 100.0
mfeat-fourier 86.28 85.78 85.23 86.48
online_shoppers 89.70 89.65 90.19 90.53
page-blocks 96.89 97.23 96.71 96.68
pc3 88.54 89.16 89.78 88.75
pendigits 99.20 99.39 99.35 99.40
rl 66.99 79.74 71.96 83.75
satimage 89.61 90.77 89.59 91.04
segment 91.49 91.40 91.37 92.77
sylvine 94.48 94.22 94.71 95.88
taiwanese_bankruptcy_prediction 96.90 96.68 96.74 96.74
waveform-5000 85.89 86.02 86.11 86.24
website_phishing 89.45 89.94 90.14 87.53
wine-quality-white 54.98 64.41 55.17 63.27
Ailerons 0.000157 0.000158 0.000155 0.000158
CookbookReviews 1.560657 1.491684 1.520183 1.486840
IEEE80211aa-GATS 0.030513 0.027608 0.029589 0.026295
Large-scale_Wave_Energy_Farm_Sydney_49 4504.285 4369.826 4369.902 47399.63
Superconductivty 10.53080 10.39726 10.64536 10.37837
archive2 388.8410 379.5506 390.6574 361.8616
bank8FM 0.028792 0.029047 0.028635 0.028826
communities_and_crime 0.143486 0.139045 0.136859 0.138984
fried 1.013058 1.013431 1.008568 1.010952
healthcare_insurance_expenses 4935.203 4743.384 4615.580 4698.968
house_16H_reg 31647.43 31333.32 30671.63 31017.80
kin8nm 0.068684 0.070110 0.067355 0.069678
mv 0.027143 0.019089 0.030948 0.027385
Table 10: Performance of TabPFN v2 and the extracted feature embeddings across 29 classification datasets. The table shows the average classification accuracy for each dataset when using different layers (Layer 6, Layer 9, Layer 12) of the transformer as feature embeddings, as well as the “combined” approach, where embeddings from up to three selected layers are concatenated. The “selected layers” column indicates the specific layers chosen for each dataset based on validation set performance. “Vanilla” refers to the embeddings extracted using the vanilla strategy, which utilizes only the 12th layer of the transformer.
PFN-v2 Vanilla layer-6 layer-9 layer-12 combined selected layers
FOREX_audchf-day-High 77.38 72.48 68.39 73.57 74.11 77.11 (5, 9, 11)
taiwanese_bankruptcy_prediction 96.99 96.70 97.14 96.77 97.07 97.14 (6)
rl 85.51 85.41 66.90 69.52 86.72 87.53 (11, 12)
pc3 89.46 89.78 90.10 88.82 88.82 88.82 (8)
eye_movements_bin 61.83 62.75 59.72 59.40 62.16 62.16 (6, 9, 12)
BNG(breast-w) 98.43 98.51 98.34 98.46 98.67 98.51 (6, 9)
FOREX_cadjpy-hour-High 69.53 70.77 62.12 64.87 70.66 70.88 (4, 5, 6)
dis 99.34 99.07 98.41 98.28 99.34 99.47 (4, 5, 6)
sylvine 97.46 97.27 92.49 93.95 97.27 96.49 (1, 11)
BNG(tic-tac-toe) 78.04 78.75 73.96 73.71 78.75 79.03 (5, 10, 12)
online_shoppers 90.59 90.11 90.02 90.11 90.63 90.02 (8)
Cardiovascular-Disease-dataset 72.84 73.08 72.96 73.06 73.14 73.09 (5, 8, 12)
credit 78.04 77.06 77.62 77.80 77.95 77.59 (4, 6, 9)
FOREX_audsgd-hour-High 67.26 69.30 57.24 61.06 69.62 70.41 (7, 10, 12)
waveform-5000 86.00 85.30 85.60 85.60 86.40 86.90 (1, 6, 11)
jungle_chess 85.65 85.30 78.55 80.44 86.66 86.85 (10, 11, 12)
BNG(cmc) 57.40 57.78 56.19 56.72 57.72 57.88 (9, 10, 12)
page-blocks 97.35 97.17 96.07 96.71 97.17 97.35 (6, 7, 12)
segment 93.07 93.29 91.99 88.10 93.51 92.64 (1, 12)
website_phishing 90.77 88.93 85.98 87.08 91.88 91.88 (7, 10)
baseball 93.66 94.03 93.28 94.03 93.66 95.15 (10, 11)
pendigits 99.50 99.36 92.81 93.04 99.41 99.45 (3, 4, 12)
Gender_Gap_in_Spanish_WP 60.84 55.79 59.68 60.32 60.53 60.84 (2, 12)
wine-quality-white 62.35 62.04 54.08 55.31 63.57 64.39 (8, 11, 12)
satimage 91.21 92.07 88.72 88.65 91.91 91.91 (8, 11, 12)
mfeat-fourier 90.00 89.75 77.75 82.25 89.50 89.50 (2, 7, 12)
VulNoneVul 98.95 98.95 98.95 98.95 98.95 98.95 (1)
law-school-admission-bianry 100.0 100.0 100.0 100.0 100.0 100.0 (6)
KDD 80.34 78.95 79.34 78.35 81.23 79.94 (1, 8, 10)
Appendix D Influence of Key Modules
We investigate the influence of two key components in TabPFN v2, i.e., the feature engineering that pre-process the raw features of a given tabular dataset and the post-hoc ensembleing.
Feature engineering. TabPFN v2 pre-processes the features of a given tabular datset with various strategies, such as quantile, category shuffling, SVD, and power transform. Specifically, we examine the effects of adding fingerprint features (add_fingerprint_feature) and polynomial features (polynomial_features) to the raw tabular data. The results indicate that TabPFN v2 performs well even without the use of these engineered features, suggesting that, for the benchmark datasets of Ye et al. (2024b), these specific feature engineering techniques do not provide a significant improvement. This finding highlights the robustness of TabPFN v2 and its ability to handle raw features effectively, without the need for extensive pre-processing or feature construction. We show the influence of this step in Figure 9.
Model ensemble. Post hoc ensembling (PHE) involves applying TabPFN v2 to the datasets multiple times with different perturbations and aggregating the predictions of these base models at different temperatures. We show the change of performance of TabPFN v2 w.r.t. the number of ensemble numbers (i.e., the number of base models) in Figure 10. On the benchmark of Ye et al. (2024b), we observe that, overall, ensemble methods improve performance, with larger ensemble sizes yielding better results. However, we also note that even without ensembling, TabPFN v2 performs exceptionally well, and the relative performance gain from ensembling is limited. This suggests that while ensembling can provide further improvements, the base TabPFN v2 model is already highly effective on its own. The equivariant property described in (Arbel et al., 2025) provides insight into this phenomenon. Since TabPFN v2 introduces random tokens to handle heterogeneous features, the model becomes less sensitive to the arbitrary ordering of features, effectively enforcing equivariance in this aspect. As a result, the benefits of ensembling through feature order permutations are less pronounced compared to TabPFN v1.

### v2

@@ Appendix E Detailed Results
Table 11: Performance of various models on 18 large-scale datasets. The results show the mean accuracy/RMSE of different models, including MLP, Logistic Regression/Linear Regression (LR), KNN, XGBoost (XGB), Random Forest (RForest), CatBoost (CatB), ModernNCA (MNCA), RealMLP, and various versions of TabPFN v2: original TabPFN v2 (PFNv2), TabPFN v2 with K-means (PFNv2-K), TabPFN v2 with Bagging (PFNv2-B), PFNv2* (TabPFNv2*), PFNv2-DT (TabPFN-DT), and PFNv2-DF (TabPFN-DF).
Dataset MLP LR KNN XGB RForest CatB MNCA RealMLP PFNv2 PFNv2-K PFNv2-B PFNv2* PFNv2-DT PFNv2-DF
BNG(credit-a) 90.07 85.98 87.41 90.21 89.25 91.13 89.98 90.91 89.55 89.01 89.66 89.89 90.45 90.43
CDC_Indicators 86.79 86.55 86.39 86.76 86.60 86.78 86.76 86.76 86.65 86.68 86.69 86.74 86.75 86.70
Higgs 75.53 64.29 65.16 73.33 71.87 74.81 73.28 75.36 71.64 71.56 72.01 72.13 73.53 73.62
Smoking_signal 73.90 72.53 72.36 73.87 73.08 73.99 73.63 74.00 73.47 73.37 73.55 73.69 73.74 73.84
nomao 96.19 94.59 95.20 96.92 96.07 97.03 96.68 96.37 96.08 96.29 96.12 96.18 96.75 96.34
sf-police-incidents 87.84 87.84 85.87 87.68 87.84 87.87 - 87.84 87.84 87.84 87.84 87.84 87.84 87.84
Data_Crowdfunding 96.48 67.04 93.70 96.89 95.29 96.81 96.53 96.71 94.59 91.81 94.96 95.07 96.90 96.83
Fashion-MNIST 89.54 85.69 86.00 90.03 86.57 90.24 89.36 90.25 68.40 82.82 83.89 86.26 78.91 78.82
covertype 94.01 72.54 92.76 96.30 78.30 90.77 97.31 97.38 83.54 82.95 84.16 86.85 97.38 97.44
jannis 71.99 64.60 65.67 71.83 69.19 72.26 72.57 73.00 70.24 70.26 70.59 71.31 72.57 72.50
poker-hand 99.99 50.12 54.01 99.51 64.63 97.69 76.31 99.88 41.97 38.86 36.80 54.12 91.13 92.33
volkert 69.85 58.75 67.41 69.74 62.71 70.88 77.18 73.76 62.82 62.15 62.81 64.84 68.66 67.76
Airlines_DepDelay (×$\times 10^{1}$) 2.905 2.933 3.170 2.891 2.907 2.881 - 2.482 2.937 2.933 2.937 2.915 2.900 2.897
Wave_Energy_Farm (×$\times 10^{3}$) 8.199 13.19 32.29 6.917 7.294 7.173 6.148 59.05 7.214 8.375 7.063 10.506 6.616 6.785
UJIndoorLoc (×$\times 10^{0}$) 9.958 $\infty$ 9.004 10.47 23.19 9.139 5.990 65.34 66.49 7.825 7.435 9.538 14.404 7.472
blogfeedback (×$\times 10^{1}$ ) 2.387 $\infty$ 2.410 2.093 2.026 2.044 1.953 2.105 3.073 2.687 2.700 2.014 1.914 1.944
microsoft (×$\times 10^{-1}$ ) 7.577 7.782 8.284 7.514 7.566 7.453 7.573 5.077 7.735 7.981 7.720 7.612 7.944 7.728
yahoo (×$\times 10^{-1}$ ) 7.692 7.997 8.504 7.629 - 7.514 - 5.671 8.148 8.332 8.132 7.961 16.409 8.069
Table 12: Performance of TabPFN v2 and the extracted feature embeddings across 29 classification datasets. The table shows the average classification accuracy for each dataset when using different layers (Layer 6, Layer 9, Layer 12) of the transformer as feature embeddings, as well as the “combined” approach, where embeddings from up to three selected layers are concatenated. The “selected layers” column indicates the specific layers chosen for each dataset based on validation set performance. “Vanilla” refers to the embeddings extracted using the vanilla strategy, which utilizes only the 12th layer of the transformer. “S” and “P” refer to unsupervised embedding extraction approaches by appending a column of dummy labels with zero values and permuting each column as labels, respectively, as described in Section C.2.
PFN-v2 Vanilla S P layer-6 layer-9 layer-12 combined selected layers
FOREX_audchf-day-High 77.38 50.68 56.95 69.48 68.39 73.57 74.11 77.11 (5, 9, 11)
taiwanese_bankruptcy_prediction 96.99 56.45 96.77 95.75 97.14 96.77 97.07 97.14 (6)
rl 85.51 50.00 60.56 70.82 66.90 69.52 86.72 87.53 (11, 12)
pc3 89.46 10.22 89.78 86.90 90.10 88.82 88.82 88.82 (8)
eye_movements_bin 61.83 50.00 55.12 57.95 59.72 59.40 62.16 62.16 (6, 9, 12)
BNG(breast-w) 98.43 69.51 97.60 98.51 98.34 98.46 98.67 98.51 (6, 9)
FOREX_cadjpy-hour-High 69.53 51.79 66.55 71.12 62.12 64.87 70.66 70.88 (4, 5, 6)
dis 99.34 85.43 98.41 98.54 98.41 98.28 99.34 99.47 (4, 5, 6)
sylvine 97.46 85.66 72.78 95.71 92.49 93.95 97.27 96.49 (1, 11)
BNG(tic-tac-toe) 78.04 34.71 71.41 73.79 73.96 73.71 78.75 79.03 (5, 10, 12)
online_shoppers 90.59 84.51 85.93 89.46 90.02 90.11 90.63 90.02 (8)
Cardiovascular-Disease-dataset 72.84 50.86 68.73 72.60 72.96 73.06 73.14 73.09 (5, 8, 12)
credit 78.04 62.31 75.86 78.31 77.62 77.80 77.95 77.59 (4, 6, 9)
FOREX_audsgd-hour-High 67.26 51.48 65.49 70.14 57.24 61.06 69.62 70.41 (7, 10, 12)
waveform-5000 86.00 80.60 55.70 87.10 85.60 85.60 86.40 86.90 (1, 6, 11)
jungle_chess 85.65 39.60 64.12 72.14 78.55 80.44 86.66 86.85 (10, 11, 12)
BNG(cmc) 57.40 42.62 52.48 55.16 56.19 56.72 57.72 57.88 (9, 10, 12)
page-blocks 97.35 94.25 95.43 96.35 96.07 96.71 97.17 97.35 (6, 7, 12)
segment 93.07 72.29 69.26 87.23 91.99 88.10 93.51 92.64 (1, 12)
website_phishing 90.77 36.90 82.66 90.04 85.98 87.08 91.88 91.88 (7, 10)
baseball 93.66 78.73 92.54 92.16 93.28 94.03 93.66 95.15 (10, 11)
pendigits 99.50 59.75 72.40 98.18 92.81 93.04 99.41 99.45 (3, 4, 12)
Gender_Gap_in_Spanish_WP 60.84 33.68 59.47 60.84 59.68 60.32 60.53 60.84 (2, 12)
wine-quality-white 62.35 10.51 49.29 55.10 54.08 55.31 63.57 64.39 (8, 11, 12)
satimage 91.21 82.04 84.99 89.19 88.72 88.65 91.91 91.91 (8, 11, 12)
mfeat-fourier 90.00 55.50 46.75 85.75 77.75 82.25 89.50 89.50 (2, 7, 12)
VulNoneVul 98.95 1.05 98.95 98.33 98.95 98.95 98.95 98.95 (1)
law-school-admission-bianry 100.0 99.83 79.76 98.82 100.0 100.0 100.0 100.0 (6)
KDD 80.34 78.45 62.36 76.76 79.34 78.35 81.23 79.94 (1, 8, 10)